When driven by a 120Hz vibrator, a string has transverse waves of 30cm wave length traveling along it. What is the speed of the waves on the string and if the tension is 1.2N, what is the mass of 50cm of the string?
Answer #1

f = 120 Hz, d=30cm = 0.3m. v=ff V=(0-3) (120) d = m/s speed of waves (v) - 36 on string F=102 N M = 9.259610 kgfm m = (9.259x